OpenStudy (anonymous):

for questions 3 and 4, what values for theta(0 less than or equal to. theta. less than or equal to 2pi) 3. 3sin(theta) = sin(theta) -1 A. pi/3, 5pi/3 B. 2pi/3, 4pi/3 C. pi/6, 3pi/6 D. 7pi/6, 11pi/6 4. tan^2(theta) = -3/2 sec(theta) A. pi/2, 5pi/2 B. 3pi/4, 5pi/4 C. pi/3, 5pi/3 D. 2pi/3, 4pi/3

OpenStudy (rajee_sam):

Just substitute the values given and find the value of the equation and se which one satisfies the condition.

OpenStudy (anonymous):

answered yesterday

OpenStudy (anonymous):

im sorry but i have absolutely no idea how to do these problems..

OpenStudy (anonymous):

\[3\sin \theta=\sin \theta-1\] \[3\sin \theta-\sin \theta=-1\] \[\sin \theta=\frac{ -1 }{2 }=-\sin \frac{ \pi }{ 6 }=\sin \left( \pi+\frac{ \pi }{6 } \right),\sin \left( 2\pi-\frac{ \pi }{ 6 } \right)\] \[\theta=\frac{ 7\pi }{ 6 },\frac{ 11\pi }{ 6 }\]

OpenStudy (anonymous):

can you help me with the other one too?

OpenStudy (anonymous):

And thanks!